# Evaluating Concurrent Robustness of Language Models Across Diverse Challenge Sets

### v2

@@ 6 Conclusion and Future Works
We demonstrate that input perturbation poses difficulties for LMs at all scales. While fine-tuned models on a single challenge set can produce robust models, their generalizability to unfamiliar perturbations remains questionable. This motivates the problem of multi-set inoculation, aiming to train a singular model resilient to a myriad of distinct perturbations. We introduce a comprehensive framework to systematically evaluate LM robustness against multiple input perturbations. In addition, we propose three strategies to fine-tune the model on multiple challenge sets. Our results underscore the superiority of mixed fine-tuning in training robust models. Furthermore, we expend our framework to LLMs, leveraging a COT prompting enriched with exemplar demonstrations.
@@ Acknowledgements
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-19-1-2620. We extend our gratitude to the annotators who verified our flowcharts and corresponding question answer pairs. Lastly, we extend our appreciation to the reviewing team for their insightful comments.
@@ A.3 Implementation Details
For LLMs: We used GPT-3.5 model and LLaMA-2 models for our experiments. GPT-3.5 has been used with a temperature setting of 0.3 (to preserve reproducibility) and 1000 maximum new tokens. LLaMA-2 model has been used after quantization with QLoRA Dettmers et al. (2023), with nf4 4-bit quantization. Double quantization has been employed and torch.bfloat16 has been used for computations during the quantization. For API calls on GPT-3.5, we have used CPU only. The cost for fine-tuning is: $0.008 for training,$0.012 for usage input, $0.016 for usage output for 1k tokens. The cost for prompting is $0.008 for 1k tokens. The number of examples are highlighted in the Section 3 and 6.

### v3

@@ 6 Conclusion and Future Works
We demonstrate that input perturbation poses difficulties for LMs at all scales. While fine-tuned models on a single challenge set can produce robust models, their generalizability to unfamiliar perturbations remains questionable. This motivates the problem of multi-set inoculation, aiming to train a singular model resilient to a myriad of distinct perturbations. We introduce a comprehensive framework to systematically evaluate LM robustness against multiple input perturbations. Additionally, we propose three strategies to fine-tune the model on multiple challenge/pertubations sets. Our results underscore the superiority of mixed fine-tuning in training robust models. Furthermore, we expand our framework to LLMs, leveraging a COT prompting enriched with exemplar demonstrations.
@@ Acknowledgements
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-23-1-2364. We extend our gratitude to the annotators who verified our flowcharts and corresponding question answer pairs. Lastly, we extend our appreciation to the reviewing team for their insightful comments.
@@ A.3 Implementation Details
For LLMs: We used GPT-3.5 model and LLaMA-2 models for our experiments. GPT-3.5 has been used with a temperature setting of 0.3 (to preserve reproducibility) and 1000 maximum new tokens. LLaMA-2 model has been used after quantization with QLoRA Dettmers et al. (2023), with nf4 4-bit quantization. Double quantization has been employed and torch.bfloat16 has been used for computations during the quantization. For API calls on GPT-3.5, we have used CPU only. The cost for fine-tuning is: $0.008 for training,$0.012 for usage input, $0.016 for usage output for 1k tokens. The cost for prompting is $0.008 for 1k tokens. The number of examples are highlighted in the Section 3 and 4.